OpenStudy (stephanos100):

PLS HELP!!! TOKENS!! Find the rectangular coordinates of the point with the polar coordinates (-3, 5 pi divided by 3)

OpenStudy (anonymous):

\[(-3, \frac{ 5\pi }{ 3 })=(r,\theta) , x=rcos \theta , y=rsin \theta\]

OpenStudy (anonymous):

Use the equations to find (x,y)

OpenStudy (stephanos100):

So it would be (-3/2, 3 square root 3/2)

OpenStudy (stephanos100):

@Wesran

OpenStudy (anonymous):

(-3/2, -3*sin(5*Pi/3))